OpenStudy (anonymous):

Given the function g(x) = 8x − 2, compare and contrast g(−2) and g(4). Choose the statement that is true concerning these two values. The value of g(−2) is larger than the value of g(4). The value of g(−2) is the same as the value of g(4). The value of g(−2) is smaller than the value of g(4). The values of g(−2) and g(4) cannot be compared.

OpenStudy (anonymous):

find g(-2) g(-2)=8(-2)-2=-16-2=-18 g(4)=8(4)-2=32-2=30 so true is the following one The value of g(−2) is smaller than the value of g(4).

OpenStudy (anonymous):

Thanks for the help :)